OpenStudy (anonymous):

What is the the solution of log3x-7 25=2 A- x=-2 B-x=2 C-x=3 D-x=4

OpenStudy (solomonzelman):

I don't understand what is the base and the log in your log ?

OpenStudy (solomonzelman):

Is it \[\huge\color{blue}{ \log _{3x-7}25=2 }\]

OpenStudy (solomonzelman):

That??? @linds ?

OpenStudy (solomonzelman):

I'll assume that's the question if you don't say anything.

OpenStudy (anonymous):

Yes it is. I am new to open study and trying to figure out how to work it, sorry

OpenStudy (solomonzelman):

\[\color{blue}{ 2=~~2~\log_{5}(5) =\log_{5}(5^2) =\log_{5}(25) }\] so lets rewrite it, \[\huge\color{blue}{ \log _{\color{red} { 3x-7 }}(25)~=~\log_{\color{red} { 5 }}(25) }\] \[\huge\color{red}{ 3x-7~~=~~5 }\] \[\huge\color{red}{ 3x=12,~~~~~~~~~~x=4 }\]

OpenStudy (solomonzelman):

It's okay. Just reply accordingly when you are asked to.

OpenStudy (solomonzelman):

Are you clear with how I solved for x ?

OpenStudy (anonymous):

Yes I am, thanks!

OpenStudy (solomonzelman):

You welcome! If you have a question, ask don't be shy!